OpenStudy (anonymous):

Find the fourth roots of the complex number z1= 1+ sqrt3 I Part 1:write z1 in polar form 2(cos60+isin60) Part 2: find the modulus of the roots of z1 I got 2 Part 3: find the four angles that define the fourth roots of the number z1 Part 4: what are the fourth roots of z1= sqrt3+1 i

OpenStudy (anonymous):

am I correct on part 1&2 and I don't know how to do 3&4

OpenStudy (irishboy123):

.

OpenStudy (anonymous):

@IrishBoy123 ??

OpenStudy (anonymous):

If \(z\) has angle \(\theta\), then the \(n\)th roots \(z^{1/n}\) will follow a pattern of \(\dfrac{\theta+2k\pi}{n}\), where \(k=0,1,\ldots,n-1\).

OpenStudy (anonymous):

@SithsAndGiggles so if the angles i find are 15 and 60 the 4th roots would be \[60+2k \div4\]?

OpenStudy (anonymous):

Idk i just have no idea how to do this

OpenStudy (anonymous):

You found that \(\theta=60^\circ\), right? In radians, that's \(\dfrac{\pi}{3}\). Take \(k=0\). Then the angle of the first (principal) fourth root is \(\dfrac{\dfrac{\pi}{3}+2\pi\times0}{4}=\dfrac{\pi}{12}\), which in degrees is \(15^\circ\). Now take \(k=1\). This gives you an angle of \(\dfrac{\dfrac{\pi}{3}+2\pi\times1}{4}=\dfrac{7\pi}{12}\), or \(105^\circ\). Continue the pattern for \(k=2\) and \(k=3\).